OpenStudy (anonymous):

Emily sets up a lemonade stand. She has set-up costs of $12.00 and each cup of lemonade costs her $0.15 to make. She sells each cup of lemonade for $0.75. Find the sum of the answers to the following questions. What is the total cost, including the set-up, for her to make 100 cups of lemonade? What is her profit (money earned minus total cost) if she sells 100 cups of lemonade? What is the number of cups that she must sell to break even (that is, to have a profit of $0)? Is it possible for her to make a profit of exactly $17.00? If it is possible, answer 10. If not, answer 5. Emily sets up a lemonade stand. She has set-up costs of $12.00 and each cup of lemonade costs her $0.15 to make. She sells each cup of lemonade for $0.75. Find the sum of the answers to the following questions. What is the total cost, including the set-up, for her to make 100 cups of lemonade? What is her profit (money earned minus total cost) if she sells 100 cups of lemonade? What is the number of cups that she must sell to break even (that is, to have a profit of $0)? Is it possible for her to make a profit of exactly $17.00? If it is possible, answer 10. If not, answer 5. @Mathematics

OpenStudy (anonymous):

MEDALS WILL BE AWARDED! :D

OpenStudy (anonymous):

Total cost = 12 + .15*100 = $27 Profit = (.75*100)-27 = 75-27= $48 Even break = 20 cups Not possible to make exactly $17 as after even break, every extra she makes gives her $0.60 profit, 17 is not divisible by 0.6 and so is not possible.